# Exposía: Academic Writing Assessment of Exposés and Peer Feedback

### v1

@@ 6.1 Experimental Setup
We use Exposía for benchmarking performance on two tasks: (i) criterion-based exposé assessment and (ii) criterion-based review assessment. Models predict scores for individual criteria.
@@
LLM prompting We assess LLM performance in a zero-shot setting. We evaluate four state-of-the-art open-weight instruction-tuned LLMs with different architectures and sizes (Appendix H). For both tasks, we use a unified prompting strategy and compare two prompting variants: (i) single-criterion prompts, which present and score one criterion at a time, and (ii) combined prompts, which list all criteria for the respective task in a single call. Cross-criterion synergy could potentially benefit the model, with a lower execution cost. Appendix I provides more detail about the prompts.
Human Qwen3 (80B) Llama 3.3 (70B) GPT OSS (120B) GPT OSS (20B)
Exp. #Crit. $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$ $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$ $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$ $\overline{A}_{\text{all}}$ $\overline{A}_{\text{ind}}$
Exposé criterion-based scoring (over 31 criteria)
0 3 0.89   0.91   0.85   0.90   0.90   0.75   0.72   0.73   0.81
1 20 0.79   0.76   0.79   0.79   0.78   0.81   0.78   0.82   0.78
2 8 0.64   0.57   0.62   0.63   0.62   0.68   0.69   0.68   0.69
All 31 0.76   0.72   0.75   0.76   0.75   0.77   0.75   0.77   0.76
Review criterion-based scoring (over 22 criteria)
1 17 0.89   0.89   0.85   0.91   0.84   0.82   0.75   0.77   0.71
2 5 0.86   0.87   0.82   0.86   0.77   0.67   0.63   0.64   0.63
All 22 0.88   0.89   0.84   0.90   0.82   0.78   0.72   0.74   0.69
Table 4: Quadratic weighted agreement (QWA) for the two tasks, grouped by expertise level (3 and 2 expertise levels for exposé and review criteria, respectively). Exp.=expertise level; #Crit.=number of criteria in the subset. Prompting variants: all (combined) and ind. (individual). Each LLM column shows human-LLM QWA averaged over both human raters and all criteria. Human column shows agreement between the human raters.
Evaluation Following studies on criterion-based scoring Pack et al. (2024); Huang and Wilson (2025), we evaluate performance, by comparing scores predicted with an LLM with a score assigned by a human rater, using QWA (Section 5.2). Results are reported on 55 exposés drafts and 110 student reviews. Unless otherwise specified, we compute QWA against Group 1 and Group 2 raters and average the results. For each exposé and review, a corresponding model predicts 31 and 22 scores, respectively,9995 exposé and 2 review criteria are excluded (Appendix H). for a total of 1705 exposé criteria scores and 2420 review criteria scores.
6.2 Key Results
Table 4 shows main results on exposé and review assessment (averaged over all criteria – rows All).
Single-criterion vs. combined prompting Across tasks and models, combined prompting yields higher agreement, while being more computationally efficient. The only exception is Qwen3 model on exposé scoring. Tables 22 and 23 tabulate criterion- and rubric-level results.
Performance by expertise level We further group criteria by expertise level: three levels for exposé, and two for review, based on the degree of domain knowledge required for assessment (Appendix D). Figure 4 demonstrates that exposé criteria requiring more in-domain knowledge display more variability, and, importantly, this aligns with human raters. Appendix Figure 7 reveals that for review scoring, the difference is not as dramatic, suggesting that review assessment does not require as much domain knowledge. See Appendix J for further discussion.
6.3 Additional Analyses
Systematic group bias
We observe a tendency for Group 1 raters to assign higher scores (Appendix G). This matters for interpreting LLM–human agreement: LLM agreement with Group 1 is often higher than agreement with Group 2, consistent with LLMs being more generous (Figure 5). We do not observe similar behavior for review scoring (Appendix Figure 8). This is consistent with the more dramatic differences on expertise level for exposé scoring and suggests that review assessment does not require as much in-domain expertise.
@@ 6.3 Additional Analyses
Scalability and classroom deployment Combined prompting is both more effective and more efficient computationally (Appendix J and K). This is important because student activity is strongly deadline-driven (Figure 6). Thus, combined prompting is recommended for classroom deployment.
7 Discussion
Interpreting model performance relative to human raters Human agreement varies substantially across criteria and expertise levels, indicating that some aspects of academic writing and feedback are inherently more difficult to assess consistently. Within this context, LLMs achieve agreement levels that approach the consensus of human agreement. Notably, performance degrades on expert-level criteria, mirroring the same dimensions where human agreement is weakest.
LLM alignment with human groups The tendency of LLMs to align more closely with Group 1 raters suggests that LLMs adopt a more generous grading regime Flodén (2024), likely to originate from the positive bias in LLMs Jain et al. (2025).
Potential tasks enabled by Exposía Beyond the tasks studied here, Exposía enables a range of research directions. These include modeling draft-feedback-revision links across documents, with numerous applications, such as feedback text generation. More broadly, the dataset supports studying how feedback mediates learning and writing development in higher-education settings.
@@ 8 Conclusion
We introduce a novel task definition for exposé and review scoring in an educational context and present the first public dataset, Exposía, of student research exposés and feedback, accompanied by human assessment scores that are based on the fine-grained grading schema we develop.
By integrating drafting, feedback, and revision into a single resource, Exposía enables pedagogically grounded research on writing and feedback in higher education and provides a unified framework that allows for a joint study of feedback and exposé writing quality. Although the dataset is collected in the Computer Science domain, the task formulation and the assessment framework are general enough to inform the development of assessment models in other areas.
We establish baseline performance with state-of-the-art LLMs: models achieve high agreement with human raters on criteria that do not require in-domain expertise, while higher expert-level criteria remain challenging. Our findings provide insight and recommendations for classroom deployment.
@@
This study has several limitations. First, Exposía is shaped by the course consent protocol: since only consented materials are included, not all exposés retain the complete set of three reviews. For the same reason, some of the reviews in Exposía are orphaned, as the author of the corresponding exposé did not provide consent (Section 5.1). As such, the consent structure reduces the number of complete draft–review–revision instances in the released data. We expect these constraints to diminish as the course is offered repeatedly and the dataset continues to grow (Section 3). Although the current size of Exposía is modest compared to some other NLP datasets, it is substantial for the educational domain, particularly given its complexity and the richness of its annotations. We discuss the time investment to create the various Exposía components (feedback and criterion-based scores on exposés and reviews) in Appendix F.3.
Second, although the course covers three areas – Human–Computer Interaction, Natural Language Processing, and Data Mining – all topics remain within Computer Science and AI. However, we believe that the task formulation and the assessment framework are general enough and can inform the development of assessment models in other areas.
Third, our model comparison prioritizes reproducibility and transparency in an educational setting and therefore focuses on openly available models. As a result, we do not evaluate proprietary closed-source LLMs, which may behave differently and limit the scope of our model comparison.
@@
The data collection was approved by the University’s Institutional Review Board and Data Protection Office under IRB proposal EK 77/2024. Data was collected based on informed consent (all instructors and 55 students), following a three-tier structure: (1) general consent for data processing within the review platform, (2) optional consent for contributing textual and highlight data, and (3) optional consent for contributing interaction data. Participation in the study was strictly voluntary and had no effect on grading. No monetary compensation was provided to students, as participation was part of the course setting; importantly, contributing research data was optional. Instructors were compensated via regular University employment contracts, following institutional pay scales and policies. The compensation was not contingent on participation in the study.
@@ Acknowledgments
This work has been funded by the German Research Foundation (DFG) as part of the PEER project (grant GU 798/28-1) and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We thank Dr. Ekaterina Kochmar (Mohamed bin Zayed University of Artificial Intelligence) and Dr. Martin Greisel (University of Augsburg) for their valuable feedback and discussions.
@@ Appendix C Additional Details on the Exposía Dataset
For users who provided explicit additional consent, we release interaction logs from the review platform, covering 45 students and 9 instructors. Each record contains an action type, a JSON-encoded data payload, a timestamp, and an anonymized user identifier. Logged events are timestamped UI interactions (Table 5); an overview of activity over the semester is shown in Figure 6 and discussed in Appendix L . Actions including navigation and visibility changes (e.g., Navigation step, PDF page show/hide, Browser Tab show/hide), scrolling and resizing (e.g., Annotation Page scroll, PDF scroll, Sidebar scroll, PDF (Browser) resize), interface controls and modal events (e.g., Sidebar click, Topbar click, Modal show/Modal hide, Modal Button click), and lightweight editing interactions (e.g., Text select, Copy text, Text paste).
@@
Appendix D Criteria and Rubrics for Exposé and Review Assessment
This appendix provides an overview of exposé and review criteria used to grade the exposé and student reviews, respectively.
@@ Appendix D Criteria and Rubrics for Exposé and Review Assessment
D.1 Exposé Criteria
@@
D.2 Review Criteria
Table LABEL:tab:review_rubric lists all the 24 review assessment criteria grouped into eight high-level rubrics. These criteria define the scoring rules applied to each submitted student review.
@@ D.2 Review Criteria
Table 6: 36 criteria used to grade student exposés. We organize the criteria into nine rubrics, by grouping together criteria that evaluate related aspects of exposé writing.
@@
Table 7: 24 criteria used to grade student reviews. We organize the criteria into eight rubrics, by grouping together criteria that evaluate related aspects of review.
@@ F.1 Exposé Revision Statistics
Table 9 reports section-wise text statistics for Motivation, Approach, and Schedule in draft vs. final versions (mean $\pm$ SD). Raw is converted to plain text with pylatexenc [^12] https://github.com/phfaist/pylatexenc; sentence/word segmentation uses spaCy. We report length/structure (#Words, #Sentences and average word length), lexical/readability (type–token ratio (TTR) and Flesch Reading Ease (FRE); Flesch, 1948 via textstat [^13] https://github.com/textstat/textstat), and draft–final similarity (cosine similarity of SentenceTransformer embeddings141414sentence-transformers/all-mpnet-base-v2 and normalized token-level Levenshtein similarity; Levenshtein, 1966).

### v2

@@ 6.1 Experimental Setup
We use Exposía to benchmark LLM performance on two tasks: (i) criterion-based exposé assessment and (ii) criterion-based assessment of student reviews.
((a)) Exposé scoring
((b)) Review scoring
Figure 4: Weighted-mean binned JSD between human and LLM score distributions (lower is better), by prompting variant. Models: Opus-4.6=Claude Opus 4.6; Gem-3.1P=Gemini 3.1 Pro; GPT-5.2=GPT 5.2; OSS-120=GPT OSS (120B); OSS-20=GPT OSS (20B); L3.3-70=Llama 3.3 (70B); Qwen-80=Qwen 3 (80B).
@@
LLM prompting We assess LLM performance in a zero-shot setting across both open-weight and closed-source models (Appendix I). For both tasks, we use a unified prompting strategy and compare two prompting variants: (i) single-criterion prompts, which present and score one criterion at a time, and (ii) combined prompts, which list all criteria for the respective task in a single call. Cross-criterion synergy could potentially benefit the model, with a lower execution cost. Appendix J provides more detail about the prompts.
Evaluation Following studies on criterion-based scoring Pack et al. (2024); Huang and Wilson (2025), we compare predicted criterion scores to human ratings using the protocol of Elangovan et al. (2025) who recommend using the binned Jensen–Shannon divergence (JSD) for capturing human–machine alignment, instead of agreement metrics such as Krippendorff, Pearson, or QWA, which are designed to account for errors in human annotation, but are not appropriate for computing human–machine agreement. For each criterion, we bin items by the median human score and aggregate the results over agreement and disagreement bins. We compute JSD between the empirical human score distribution and the model’s predicted score distribution within each bin. We report the mean binned JSD (lower is better), aggregated over criteria (we also report QWA between models and human raters for completeness). Results are reported on 55 exposé drafts and 110 student reviews. For each exposé and review, the corresponding model predicts 31 and 22 scores, respectively,1010105 exposé and 2 review criteria are excluded (Appendix I). for a total of 1705 exposé and 2420 review criterion scores.
6.2 Results
Figure 4 shows key results on exposé and review assessment (averaged over all criteria). Also shown in Table 25, and Table 17 tabulates QWA results. We emphasize the JSD results, however, especially since many of our criteria exhibit relatively high human variation (Table 2), and when the proportion of samples with uncertainty in human-assigned labels is relatively high, machine labels may superficially appear to have similar or better correlation with the human majority label compared to the human-to-human correlation. This can create the illusion that labels from automatic evaluation approximate the human majority label Elangovan et al. (2025), which is exactly what we observe when evaluating using QWA. Our key findings are as follows:
@@ 6.2 Results
Combined prompting outperforms single-criterion prompting Across tasks and models, combined prompting yields higher alignments with human raters. Combined prompting is also more efficient computationally (Appendix K and L). This is important since student activity is deadline-driven (Figure 5). Thus, combined prompting is recommended for classroom deployment.
Closed-source models outperform open-weight models Across tasks, closed-source models tend to achieve lower binned JSD than open-weight models, indicating closer alignment with human score distributions. At the same time, educational deployments impose strict privacy constraints: students may not consent to their work being transmitted to public-facing model APIs. This creates a practical need for better-aligned, locally deployable open-weight models that support privacy-preserving educational applications.
Different models are better suited for exposé and review scoring We find that Claude exhibits consistently high alignment with humans on both tasks. However, other models are less consistent. For instance, GPT 5.2 attains the best alignment on exposé scoring but is among the poorest models on review scoring, whereas Llama 3.3 exhibits the opposite behavior. This suggests that different models may be better suited to the two tasks. In Appendix K, we identify criteria on which LLMs align most closely with human raters.
Model performance is stable on review scoring but not exposé scoring Finally, we observe that all models exhibit similar alignment levels on review scoring, whereas on exposé scoring the models exhibit less stability, and the difference between the top-performing model (GPT 5.2) and the most poorly-aligned Qwen 3 is substantial.
6.3 Illustrative Examples from Exposía
Appendix H presents two examples from Exposía. The examples illustrate how an exposé draft, reviewer feedback, and a revised exposé are connected in the dataset. We briefly discuss one example that focuses on the exposé criterion “Teaser” (Research Question). This example shows how the Teaser criterion is used to assess whether a proposal states a clear research question. The relevant excerpt from the draft proposal, including an inline comment from an instructor, is below:
[…]This research addresses the critical gap in AI-driven cancer research by integrating the impact of the tumor microenvironment into data analysis workflows for personalized treatment planning. We will […] address the limitations of current AI models in incorporating TME data and explore the anticipated improvement of personalized treatment prediction. [Weakness: The goals remain broad; the main research question is not clearly stated.][…]
In the draft, the research direction is understandable, but remains too broad because it describes goals and anticipated contributions rather than a distinct research question. This is reflected in the instructor reviews and the corresponding scores assigned on the Teaser criterion (1 and 0 for each instructor, for a maximum of 2, Table 14). We show a review excerpt from the instructor who gave a 1:
[…]Having an explicit, clearly defined research question gives a lot of structure to the overall expose… the exact goals and your main focus are not entirely clear to me. […]
The final exposé (Appendix H) illustrates how the exposé was successfully revised, following the feedback, and assessed at 2 on this criterion.
Lastly, we analyze LLM-generated assessment scores for this example. We find that the models disagree in their scores. The best-performing model in Figure 4 (GPT 5.2) assigned a score of 0 with the following justification: “No explicit research question is stated.”. Several models interpreted the passage as an implicit research direction: Claude Opus assigned a score of 1 and argued that the work was described in broad terms without a clearly articulated research question. Yet other models gave a score of 2. For example, Llama 3.3 described the research question as clearly formulated.
Although the instructors disagree on the score, both identified the missing Teaser. In contrast, several LLMs, especially open-source ones, failed to detect the absence of a clearly formulated research question. This example highlights the limitations of LLM-based scoring and motivates future work on alternative approaches and improved prompting based on closer analysis of model justifications.
7 Tasks Enabled by Exposía
As discussed in Section 4, the annotations in Exposía are intentionally broad and support a wide range of tasks for scientific writing assistance and assessment, beyond those evaluated in this paper. This richness is a key strength of the dataset. [^11] Publishing a dataset with additional components that enable new tasks is consistent with other resource papers in CL venues Baumel et al. (2014); Qin et al. (2018). In particular, the dataset enables research on modeling draft–feedback–revision links across documents, with numerous applications. More broadly, the dataset supports studying how feedback mediates learning and writing development in higher-education settings. Because the corpus contains paired drafts and final exposés, it enables quantifying draft-to-final improvements on specific criteria and relating these changes to the feedback received. We find that exposés improve in response to feedback (Section 5), particularly on criteria related to in-domain content rather than form, but more fine-grained analysis remains important future work and is beyond the scope of the present paper.
Exposía supports more fine-grained tasks such as linking inline comments to specific revisions, predicting revisions from feedback, and modeling feedback properties, associated with criterion-level improvements in revised exposés. Such models can help identify types of feedback that are most effective for improving specific writing dimensions.
Automatic feedback generation is another important application, supported by Exposía. Good feedback helps improve writing and, at the same time, helps students improve their own feedback Wu and Schunn (2021); Jiang et al. (2024). This connection is well known in the area of educational research, but empirical studies are lacking due to the absence of appropriate datasets that connect writing and feedback Wu and Schunn (2021).
Finally, automated scoring of exposés supports scalable assessment for research proposal writing, while automated review scoring supports scalable assessment of student ability to provide effective peer feedback (a key practical need given the substantial human grading effort; Appendix F.3).
All of these applications directly contribute to addressing the two challenges of (1) research proposal writing and (2) providing feedback on proposal writing. Exposía is the first dataset that explicitly connects the two target skills.
@@ 8 Conclusion
We introduce a novel task definition for teaching and assessment of exposé and feedback writing skills, and present the first public dataset, Exposía, of student exposés and feedback, accompanied by human assessment scores based on the fine-grained grading schema we develop.
By integrating drafting, feedback, and revision into a single resource, Exposía enables pedagogically grounded research on writing and feedback in higher education and provides a unified framework that allows for a joint study of feedback and exposé writing quality. We also outline a range of tasks enabled by Exposía. Although the dataset is collected in the Computer Science domain, the task formulation and the assessment framework are general enough to inform the development of models in other areas. We establish baseline performance with state-of-the-art LLMs on exposé and review scoring tasks. Our resource, findings and models pave the path for the holistic study of computational support for academic writing and feedback.
@@
This study has several limitations. One limitation is that Exposía is shaped by the course consent protocol: since only consented materials are included, not all exposés retain the complete set of three reviews. For the same reason, some of the reviews in Exposía are orphaned, as the author of the corresponding exposé did not provide consent (Section 5.1). As such, the consent structure reduces the number of complete draft–review–revision instances in the released data. We expect these constraints to diminish as the course is offered repeatedly and the dataset continues to grow (Section 3). We expect 600 students to enroll annually, with an estimated 30% willing to contribute their content, recognizing the importance of data collection in a course focused on teaching important skills for writing scientific papers and for providing effective feedback on scientific writing. Note that this winter term, around 500 students already enrolled in the course featuring three new research topics, and 30% of the students provided their consent. Although the current size of Exposía is modest compared to some other NLP datasets, it is substantial for the educational domain, particularly given its complexity and the richness of its annotations. We discuss the time investment to create the various Exposía components (feedback and criterion-based scores on exposés and reviews) in Appendix F.3.
Another limitation concerns the generalizability of the dataset and the assessment criteria beyond Computer Science and AI. Although the course covers three areas – Human–Computer Interaction, Natural Language Processing, and Data Mining – all topics remain within these domains. This raises the question of the extent to which Exposía, its assessment criteria, and the experimental findings presented in this work can transfer to other disciplinary contexts. We stress that the course is designed to teach research proposal writing and peer feedback skills, which generalize across disciplines and are applicable to early-career researchers in general. Furthermore, our assessment criteria are grounded in educational assessment research that is not specific to Computer Science (Section 4) but are derived from cross-disciplinary guidance documents (Appendix E). We therefore expect the proposed framework to generalize to other disciplines, although broader empirical validation remains important future work. Importantly, because the guidelines and criteria are thoroughly documented in the Appendix (see Appendix D), they can also support data collection in other subject areas.
Additionally, we leave for future work the question of empirically validating the assessment criteria that we design for exposé and review grading. This is because validating the assessment framework would require a large-scale comparative study in which students are assessed using either the criteria developed in this work, or an alternative assessment scheme, and the outcomes of both groups would need to be compared. Such a study would take more than a year to complete and would require the involvement of several researchers. It is therefore outside the scope of the present work.
Another, related limitation is that Exposía focuses specifically on research proposal writing, which differs from other (academic) writing tasks in purpose, structure, and style (see Section 2). Similarly, exposé feedback differs from other forms of feedback studied in the NLP community, in particular peer review for academic publishing. As a result, the assessment criteria and experimental findings reported in our study may not transfer directly to other writing tasks without further validation. At the same time, we expect some aspects to generalize beyond this setting: several criteria, such as language quality and organization, overlap with related areas like automated essay scoring, even though other dimensions, including motivation, methodology, and approach, are more closely tied to the goals of proposal writing. Future work should therefore investigate how the dataset, criteria, and models can inform research on other writing tasks.
A further limitation is that open-source models do not yet perform on par with the strongest closed-source systems across the tasks considered in this work. This is especially relevant in educational settings, where student consent and privacy constraints may limit the use of external proprietary models. Future work should therefore focus on developing stronger open-source assessment models and privacy-preserving deployment strategies that better fit the practical and ethical constraints of educational data.
@@
Collecting data in an educational setting is challenging due to the high sensitivity of the data, which raises serious privacy and ethical issues Guan et al. (2023). The data collection was approved by the University’s Institutional Review Board and Data Protection Office under IRB proposal EK 77/2024.
Data was collected based on informed consent (all instructors and 55 students), following a three-tier structure: (1) general consent for data processing within the review platform, (2) optional consent for contributing textual and highlight data, and (3) optional consent for contributing interaction data. Participation in the study was strictly voluntary and had no effect on grading for the students. No monetary compensation was provided to students, as participation was part of the course setting; importantly, contributing research data was optional. Instructors were compensated via regular University employment contracts, following institutional pay scales and policies. The compensation was not contingent on participation in the study.
@@ Acknowledgments
This work has been funded by the German Research Foundation (DFG) as part of the PEER project (grant GU 798/28-1), Hessian.AI Connectom Fund (LLMentor: Expert-AI Coteaching of “Introduction to Scientific Work”) and by the European Union (ERC, InterText, 101054961). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We thank Dr. Ekaterina Kochmar (Mohamed bin Zayed University of Artificial Intelligence) and Dr. Martin Greisel (University of Augsburg) for their valuable feedback and discussions.
@@ Appendix C Additional Details on the Exposía Dataset
For users who provided explicit additional consent, we release interaction logs from the review platform, covering 45 students and 9 instructors. Each record contains an action type, a JSON-encoded data payload, a timestamp, and an anonymized user identifier. Logged events are timestamped UI interactions (Table 3); an overview of activity over the semester is shown in Figure 5 and discussed in Appendix M . Actions including navigation and visibility changes (e.g., Navigation step, PDF page show/hide, Browser Tab show/hide), scrolling and resizing (e.g., Annotation Page scroll, PDF scroll, Sidebar scroll, PDF (Browser) resize), interface controls and modal events (e.g., Sidebar click, Topbar click, Modal show/Modal hide, Modal Button click), and lightweight editing interactions (e.g., Text select, Copy text, Text paste).
@@
Appendix D Guidelines on Criteria and Rubrics for Exposé and Review Assessment
This appendix provides an overview of exposé and review criteria and the guidelines used by the instructors when grading the exposés and student reviews, respectively.
@@ Appendix D Guidelines on Criteria and Rubrics for Exposé and Review Assessment
D.1 Annotation Guidelines: Exposé Grading Criteria and Rubrics
@@
D.2 Annotation Guidelines: Review Grading Criteria and Rubrics
Table LABEL:tab:review_rubric lists all the 24 review assessment criteria grouped into eight high-level rubrics. These criteria define the scoring rules applied to each submitted student review. While actionable feedback promotes learning, ineffective or misleading reviews can have the opposite effect. To this end, inspired by Du et al. (2024), the criteria include a penalty for unhelpful feedback patterns (e.g., neglect of key elements, unsupported claims, or contradictory advice).
@@ D.2 Annotation Guidelines: Review Grading Criteria and Rubrics
Table 4: Guidelines for exposé grading. 36 criteria used to grade student exposés. We organize the criteria into nine rubrics, by grouping together criteria that evaluate related aspects of exposé writing.
@@
Table 5: Guidelines for review grading. 24 criteria used to grade student reviews. We organize the criteria into eight rubrics, by grouping together criteria that evaluate related aspects of review.
@@ F.1 Exposé Revision Statistics
Table 7 reports section-wise text statistics for Motivation, Approach, and Schedule in draft vs. final versions (mean $\pm$ SD). Raw is converted to plain text with pylatexenc [^14] https://github.com/phfaist/pylatexenc; sentence/word segmentation uses spaCy. We report length/structure (#Words, #Sentences and average word length), lexical/readability (type–token ratio (TTR) and Flesch Reading Ease (FRE); Flesch, 1948 via textstat [^15] https://github.com/textstat/textstat), and draft–final similarity (cosine similarity of SentenceTransformer embeddings161616sentence-transformers/all-mpnet-base-v2 and normalized token-level Levenshtein similarity; Levenshtein, 1966).